OpenStudy (anonymous):

simplify the sxpression: −3x^2(5x − 4x^2 − 6). ^ indicates exponets

OpenStudy (anonymous):

try to use this first ! a(b+c+d)=ab+ac+ad If you want a full answer ! I can give it to you ...But...you should learn how to do it by yourself ! :)

OpenStudy (anonymous):

can you show the steps please?

OpenStudy (anonymous):

ok ! \[-3x^2(5x-4x^2-6)=(-3x^2)(5x)+(-3x^2)(-4x^2)+(-3x^2)(-6)\] Can you compute this ?! :)

OpenStudy (anonymous):

\[-15x^3+12x^4+18x^2\]

OpenStudy (anonymous):

im confused..

OpenStudy (anonymous):

did you understand the first step ?!